# Assessing biomedical knowledge robustness in large language models by query-efficient sampling attacks

### v2

@@ 5.1 Two regimes of adversarial entities
@|l @ P1cm @ |P1cm @ P1.8cm @ P1.3cm @ P1.3cm @ |P1cm @ P1.8cm @ P1.3cm @ P1.3cm |@ MedQA-USMLE (ASR %) MedMCQA (ASR %)
@@
MedAlpaca-7B Drugs 20.3 22.8 / 24.5 22.8 20.5 24.9 34.2 / 27.5 42.6 36.8
Diseases 25.2 30.1 / 21.6 26.5 25.2 31.2 41.7 / 25.0 40.5 38.1
@@
Table 2: Multi-query (here budget = 8) attack success rates (ASR %) on various LLMs by perturbing the drug and disease entity mentions in MedQA-USMLE and MedMCQA datasets. The PDWS attack uses the embedding from CODER and DiscreteZOO (DZOO) attacks use either CODER or GTE-base to calculate the cosine distance. For each model, the highest ASR for a type of entity-mention question in each dataset is bolded.
5.2 Scaling characteristics from adversarial entities
Size is not all for scaling model robustness.
We compare model robustness using single-query ASR as the metric in Table 5 and Fig. 3. For the Flan-T5 series of GLMs (Chung et al., 2024) and Flan-UL2 (Tay et al., 2022), the single-query ASR is more pronounced in smaller and less performant models, indicating the improvement of robustness as the model’s size grows. However, an opposite trend of robustness is seem in BLMs. It is worth noting that although Palmyra-Med-20B (Kamble & Alshikh, 2023) has about three times the number of parameters than MedAlpaca-7B (Han et al., 2023), the larger model is much more sensitive to entity perturbation. In terms of ASR, the BLMs are comparable to the GLMs in each type of questions evaluated on. Another observation is that the ASR for models is asymmetrical under PDWS attacks. For most of the models, sampling adversarial entities with inverse-distance-proportional weights ($n<0$ in Eq. 3) induces more performance drop than sampling with distance-proportional weights ($n>0$ in Eq. 3), yet prominent counterexamples exist. The attack performance is further investigated in the sequential, multi-query setting (budget = 8) in Table 5.1, where we also compared sampling- against search-based DiscreteZOO attacks. Here, PDWS attacks in the $n<0$ regime are still largely the most effective.
@@ 5.2 Scaling characteristics from adversarial entities
Figure 4: Example scaling curves of the query budget against ASR for disease-mention questions with (a) Flan-T5-xl on MedMCQA dataset and (b) MedAlpaca model on MedQA-USMLE dataset. The curves are generated using TCES of disease names with the methods described in the legends. The DiscreteZOO attacks were run in the low-query-budget setting. Executing the random sampling (RS) attack doesn’t require a text embedding, while the other attacks were evaluated with the CODER or GTE-base as text embedding.
Sampling in specialized text embedding has favorable query scaling over blackbox gradient attacks.
A query scaling study of adversarial attacks investigates attack success with varying query budgets per input instance (Shukla et al., 2021). It is a comprehensive evaluation of attack methods, but not yet frequently adopted in most publications, partly due to the computational cost. Fig. 4 compares the query scaling of both sampling- and search-based methods for TCES. For low-query-budget attacks (Fig. 1a), each scaling curve exhibits a rapid rise phase and a plateau phase, separated by an inflection point. The key message here is that the advantage of PDWS attacks over DiscreteZOO exists with a choice of text embedding. It is greater in specialized text embeddings but is reduced or diminished in general ones thanks to the difference in neighborhood structures. We expand the analysis in the following in three directions.
  • Sampling hyperparameter controls scaling curve shape in PDWS attacks: The hyperparameter $n$ changes the inflection point of the scaling curve. When the query budget is past the inflection point of the curve, random sampling starts to surpass PDWS ($n>0$) in ASR. This is a consequence of the distribution of adversarial entities in the embedding space and the regional bias derived from the sampling weights. The examples in Fig. 4 show that the largest performance gap between the PDWS and DiscreteZOO attacks appears at around the inflection point of the respective query scaling curve.
  • Gradient quality affects query scaling: The query complexity in blackbox gradient estimation is a long-standing theoretical topic in derivative-free optimization (L’Ecuyer & Yin, 1998) and online learning (Duchi et al., 2015). Nevertheless, in application settings such as adversarial machine learning, how gradient quality affects query scaling remains less studied for text data. Gradient estimation consumes a part of the total query budget, leading to a tradeoff between the gradient quality and terminal ASR attainable in the query scaling curve. We empirically explored the query scaling behavior of the DiscreteZOO attack in Appendix E with different candidate points in the multi-point gradient estimation, which in effect alters the gradient quality. The results show that having too few candidate points (i.e. small $M$ in Eq. 6) leads to an early onset of the plateau phase, thereby limiting the highest attainable ASR within the total query budget, whereas having many candidate points can lead to a slower growth of ASR in the rapid rise phase of the query scaling curve.
@@
Figure 5: (Top) An entity substitution attack using n-acetylglucosamine as the replacement entity for metoprolol creates an adversarial distractor. (Bottom) Heatmaps of token-wise Shapley values for a question before and after the adversarial attack on option C. The model prediction changes from the correct choice of B (unperturbed) to the incorrect choice of D (adversarially perturbed).
@@ 6 Discussion and future work
The present study illustrates that distracting information introduced through QA distractors can lead to noticeable performance degradation in billion-parameter LLMs. TCES with only a single entity modification step combined with adversarial sampling methods and appropriate text embeddings is a feasible way for collecting adversarial entity data. Moreover, TCES ensures that the replacement entity appears natural to humans and is less likely to be detected using grammar check or topical filtering (Morris et al., 2020a). The setting simulates realistic scenarios where the “attacks" may be initiated by unsuspecting users, such as biomedical professionals using an LLM-based clinical decision support system (Liu et al., 2023). In the examples illustrated in Fig. 5 and Appendix E, adversarial distractors can lead to disease misdiagnosis or misprescription of medication, which are detrimental scenarios that may occur in algorithmic decision-making in biomedicine using LLMs. Further improvement on the attack performance may be achieved using learned samplers through rejection mechanisms (Narasimhan et al., 2024) or by finetuning the text embedding to improve the query efficiency of sampling. Our approach may be integrated into interactive platforms for adversarial testing in open-ended human-machine conversation (Chao et al., 2023) for online model auditing.
Future research may focus on how to make LLMs more reliable through prompt engineering (Si et al., 2023; Nori et al., 2023) or leveraging retrieval from nonparametric knowledge bases (Weikum et al., 2021; Soman et al., 2023) or text sources (Wang et al., 2023b) to improve model generalization to long-tail entities. Besides vulnerability identification, the current work motivates adversarial defense strategies based on generalized distances to establish a tiered system for validating adversarial robustness in user-facing and domain-oriented applications.
Broader Impact Statement
The scenarios considered in this work using public datasets relate to the potential real-world deployment of LLMs in biomedical education and clinical decision support. Our results show that the lack of robustness assessment or adversarial testing methods in high-stakes domains may incur serious consequences because of naturally looking adversarial examples. The design and deployment of adversarial testing should be specific to and carried out at every stage of the model development cycle, paired with tools such as adversarially robust model explainability methods for understanding model behavior.
@@ B.1 The limit of $n\rightarrow+\infty$
$e_{j}=\mathrm{argmax}\,\mathcal{H}_{\textrm{-}k,\uparrow}=\underset{e_{j}\in\mathcal{E}_{\textrm{-}k}}{\mathrm{argmax}}\,h(k,e_{j})$ (13)
is selected with certainty at the deterministic limit.
@@ B.2 The limit of $n\rightarrow-\infty$
A similar approach to B.1 can be used to construct the probability weights in the limit $n\rightarrow-\infty$. As noted in the footnote of the main text, this limit is not attained since $k\notin\mathcal{E}_{\textrm{-}k}$.
@@ Appendix E Model evaluation and attack examples
We list some examples of successful entity substitution attacks using adversarial distractors following Fig. 5 in the main text. For each example, the text box on top (the prompt formatting removed) shows the original options on the left side and the perturbed options on the right side. The entity of interest before and after perturbation is colored in red and blue, respectively. The checkmark indicates the model prediction on unperturbed and perturbed questions. Beneath the text box are the corresponding heatmaps of token-wise Shapley values (or Shapley profile) before and after the entity substitution attack. The Shapley profiles for the MedQA-USMLE dataset (Jin et al., 2021) were generated using the model Flan-T5-large and those for the MedMCQA (Pal et al., 2022) using the model Flan-T5-xxl (Chung et al., 2024). The Shapley values were calculated using the partition explainer in shap, which uses the Owen values as an approximation (Chen et al., 2023).

### v3

@@ 5.2 Scaling characteristics from adversarial entities
@|l @ P1.1cm @ |P0.9cm @ P2cm @ P1.3cm @ P1.5cm @ |P0.9cm @ P2cm @ P1.3cm @ P1.5cm |@ MedQA-USMLE (ASR %) MedMCQA (ASR %)
@@
MedAlpaca-7B Drugs 16.8 21.1 / 22.3 22.8 20.5 27.9 37.1 / 30.1 42.6 36.8
Diseases 21.7 30.3 / 24.0 26.5 25.2 29.8 43.9 / 30.8 40.5 38.1
@@
Table 2: Multi-query (here budget = 8) attack success rates (ASR %) on various LLMs by perturbing the drug and disease entity mentions in MedQA-USMLE and MedMCQA datasets. The PDWS attack uses the embedding from CODER and DiscreteZOO (DZOO) attacks use either CODER or GTE-base embedding. For each model, the highest ASR for a type of entity-mention question in each dataset is bolded.
@@
Figure 4: Example scaling curves of the query budget against ASR for disease-mention questions with (a) Flan-T5-xl model on MedMCQA dataset and (b) MedAlpaca model on MedQA-USMLE dataset. The curves are generated using TCES of disease names with the methods described in the legends. The DiscreteZOO attacks were run in the low-query-budget setting. Executing the random sampling (RS) attack doesn’t require a text embedding, while the other attacks were evaluated with the CODER or GTE-base as text embedding. In (c) and (d), query scaling of the attacks based on $B$-nearest and $B$-farthest element sampling ($B$ is query budget) for the same datasets and models as in (a) and (b) are compared with PDWS.
@@
Figure 5: (Top) An entity substitution attack using n-acetylglucosamine as the replacement entity for metoprolol creates an adversarial distractor. (Bottom) Heatmaps of token-wise Shapley values for a question before and after the adversarial attack on choice C. The model prediction changes from the correct choice of B (unperturbed) to the incorrect choice of D (adversarially perturbed).
Sampling attacks has improved query scaling over blackbox gradient attacks in specialized text embedding.
A query scaling study of adversarial attacks investigates attack success with varying query budgets per input instance (Shukla et al., 2021). It is a comprehensive evaluation of attack methods. Fig. 4 compares the query scaling of both sampling- and search-based methods for TCES. For low-query-budget attacks (Fig. 1a), each scaling curve exhibits a rapid rise phase and a plateau phase, separated by an inflection point. The key message here is that the advantage of PDWS attacks over DiscreteZOO depends on the model and the choice of text embedding. The PDWS attacks are overarchingly more effective for GLMs than DiscreteZOO, while for BLMs, the advantage is reduced. Moreover, using a specialized text embedding further boosts the advantage of PDWS attacks compared with a general text embedding due to the difference in embedding space neighborhood. We expand the analysis in the following in three directions.
  • Sampling hyperparameter controls scaling curve shape in PDWS attacks: The hyperparameter $n$ changes the inflection point of the scaling curve. When the query budget passes the inflection point of the curve, random sampling starts to surpass PDWS ($n>0$) in ASR. This is a consequence of the distribution of adversarial entities in the embedding space and the regional bias derived from the sampling weights. The examples in Fig. 4 show that the largest performance gap between the PDWS and DiscreteZOO attacks appears at around the inflection point of the respective query scaling curve.
  • Gradient quality affects query scaling in DiscreteZOO attacks: The query complexity in blackbox gradient estimation is a long-standing theoretical topic in derivative-free optimization (L’Ecuyer & Yin, 1998) and online learning (Duchi et al., 2015). Nevertheless, how gradient quality affects query scaling remains less studied for adversarial attacks employing gradient-based discrete optimization. Gradient estimation consumes a large part of the total query budget, leading to a tradeoff between the gradient quality and terminal ASR attainable (at max budget) in the query scaling curve. We empirically explored the query scaling behavior of the DiscreteZOO attack in Appendix E with different candidate points in the multi-point gradient estimation, which in effect alters the gradient quality. The results show that having too few candidate points (i.e. small $M$ in Eq. 5) leads to an early onset of the plateau phase, thereby limiting the highest attainable ASR, whereas having many candidate points can lead to a slower growth of ASR in the rapid rise phase of the query scaling curve.
  • Deterministic sampling attacks are competitive proxies of their probabilistic counterparts. Performance comparison between $B$-nearest and $B$-farthest element sampling, the deterministic and query-limited versions of PDWS (corresponding to the $n<0$ and $n>0$ regimes) are given in Fig. 4c-d for the two example cases studied for the query scaling. The results show that the deterministic sampling attacks are already competitive against DiscreteZOO, and also close in performance to PDWS in both specialized and general text embeddings. The difference is that the two types of deterministic sampling attacks examine only predefined regions in the embedding space and cannot explore the entire attack surface as PDWS would through tuning the $n$ hyperparameter.
@@ 6 Discussion and future work
The present study illustrates that TCES is a compact yet effective way to construct adversarial distractors, which leads to significant performance degradation in LLMs. Our results suggest that to minimize model queries, using a budget of up to the inflection point on the query scaling curve is the most cost-effective. The perturbed texts produced by TCES appear natural and are less likely to be detected using grammar check or topical filtering (Morris et al., 2020a). Therefore, the setting can simulate realistic scenarios where the “attacks” may be initiated by unsuspecting users, such as healthcare professionals using an LLM-based clinical decision support system (Liu et al., 2023). In the examples illustrated in Fig. 5 and Appendix E, adversarial distractors can lead to disease misdiagnosis or misprescription of medication, which are detrimental scenarios that may occur in algorithmic decision-making in biomedicine using LLMs. Further improvement on the sampling attack performance may be achieved using learned samplers through rejection mechanisms (Narasimhan et al., 2024) or by finetuning the text embedding to improve the query efficiency of sampling. Alternatively, the semantic distance used for PDWS may be replaced with a concept distance (Choi et al., 2016) to improve attack performance on general text embeddings. Our approach may be integrated into interactive platforms for adversarial data collection or online monitoring systems for open-ended human-machine conversation (Chao et al., 2023).
Future research may focus on how to make LLMs more reliable through prompt engineering (Si et al., 2023; Nori et al., 2023) or leveraging retrieval from nonparametric knowledge bases (Weikum et al., 2021; Soman et al., 2023) or text sources (Wang et al., 2023b) to improve model generalization to long-tail entities. Besides vulnerability identification, the current work motivates adversarial defense strategies based on generalized distances (La Malfa et al., 2020) to establish a tiered system for robustness assessment in user-facing and domain-oriented applications to mitigate catastrophic failures.
Acknowledgments
R.P.X. thanks N. Rethmeier for helpful technical discussions. A.J.L. acknowledges funding from NIH T32GM067547, the UCSF Discovery Fellows Program, and the Achievement Rewards for College Scientists (ARCS) Fellowship. R.A.-A. would like to acknowledge funding from Weill Neurohub.
@@ B.1 The limit of $n\rightarrow+\infty$
$e_{j}=\mathrm{argmax}\,\mathcal{H}_{\textrm{-}k,\uparrow}\equiv\underset{e_{j}\in\mathcal{E}_{\textrm{-}k}}{\mathrm{argmax}}\,h(k,e_{j})$ (12)
is selected with certainty at the deterministic limit. The equivalence ($\equiv$) is used because the two sets -$\mathcal{H}_{\textrm{-}k,\uparrow}$ and $h(k,e_{j})$ $(j=1,...,m)$ have the same elements. Eq. 12 also defines the farthest element sampling.
@@ B.2 The limit of $n\rightarrow-\infty$
A similar approach to B.1 can be used to construct the probability weights in the limit $n\rightarrow-\infty$. As noted in the footnote of the main text, this limit is not attained since $k\notin\mathcal{E}_{\textrm{-}k}$. Nevertheless, we can define the deterministic sampling of the nearest entity to $k$ in a similar vein as Eq. 12,
$e_{j}=\mathrm{argmin}\,\mathcal{H}_{\textrm{-}k,\uparrow}\equiv\underset{e_{j}\in\mathcal{E}_{\textrm{-}k}}{\mathrm{argmin}}\,h(k,e_{j}),$ (13)
which we refer to as nearest element sampling.
@@ B.3 The case of $n=0$
B.4 Query-limited deterministic sampling
Given a query budget $B$, construct the ordered sets -$\mathcal{H}_{\textrm{-}k,\uparrow}^{j,m}=\{h_{j,\uparrow},...,h_{m,\uparrow}\}$ ($j=1,...,B$), and assuming $B\leq m$, we know that -$\mathcal{H}_{\textrm{-}k,\uparrow}^{j,m}\subseteq\mathcal{H}_{\textrm{-}k,\uparrow}\equiv\mathcal{H}_{\textrm{-}k,\uparrow}^{1,m}$ from the definition of -$\mathcal{H}_{\textrm{-}k,\uparrow}$ in Appendix B.1. We use $B$-nearest element sampling to refer to the deterministic sampling carried out sequentially where at the $j$th ($j=1,...,B$) query, the sampled entity $e_{j}=\mathrm{argmin}\,\mathcal{H}_{\textrm{-}k,\uparrow}^{j,m}$. Similarly, we can construct the ordered sets -$\mathcal{H}_{\textrm{-}k,\uparrow}^{1,m-j+1}=\{h_{1,\uparrow},...,h_{m-j+1,\uparrow}\}$ ($j=1,...,B$). We use $B$-farthest element sampling to refer to the sequential deterministic sampling where at the $j$th ($j=1,...,B$) query, the sampled entity $e_{j}=\mathrm{argmax}\,\mathcal{H}_{\textrm{-}k,\uparrow}^{1,m-j+1}$. When $B=1$, the query-limited versions reduces to the farthest and nearest element sampling described earlier in Appendix B with only a single query.
@@ Appendix E Model evaluation and attack examples
We list some examples of successful entity substitution attacks using adversarial distractors following Fig. 5 in the main text. For each example, the text box on top (the prompt formatting removed) shows the original choices on the left side and the perturbed choices on the right side. The entity of interest before and after perturbation is colored in red and blue, respectively. The checkmark indicates the model prediction on unperturbed and perturbed questions. Beneath the text box are the corresponding heatmaps of token-wise Shapley values (or Shapley profile) before and after the entity substitution attack. The Shapley profiles for the MedQA-USMLE dataset (Jin et al., 2021) were generated using the model Flan-T5-large and those for the MedMCQA (Pal et al., 2022) using the model Flan-T5-xxl (Chung et al., 2024). The Shapley values were calculated using the partition explainer in shap, which uses the Owen values as an approximation (Chen et al., 2023).